A rectangle has an area of 24cm2. What might the perimeter be?
Answer by shailendrapatel220@gmail.com(1)

(22/7)*r^2=24
r^2=(24*7)/22
r=sqrt(84/11)
r=sqrt7.7
r=2.7
perimeter=2*3.14*2.7
=17.27cm